# Smart Contract Fuzzing Towards Profitable Vulnerabilities

### v1

@@ Abstract.
DISCLAIMER: The paper has been accepted by FSE’25 and this is the original submission version without revisions requested from peer reviews. We will update camera-ready version shortly.
@@
To address these gaps, we have developed Verite, a profit-centric smart contract fuzzing framework that not only effectively detects those profitable vulnerabilities but also maximizes the exploited profits. Verite has three key features: ① DeFi action-based mutators for boosting the exploration of transactions with different fund flows; ② potentially profitable candidates identification criteria, which checks whether the input has caused abnormal fund flow properties during testing; ③ a gradient descent-based profit maximization strategy for these identified candidates. Verite is fully developed from scratch and evaluated on a dataset consisting of 61 exploited real-world DeFi projects with an average of over 1.1 million dollars loss. The results show that Verite can automatically extract more than 18 million dollars in total and is significantly better than state-of-the-art fuzzer ItyFuzz in both detection (29/9) and exploitation (58 times more profits gained on average). Remarkbly, in 12 targets, it gains more profits than real-world attacking exploits (1.01 to 11.45 times more). Verite is also applied by auditors in contract auditing, where 6 (5 high severity) zero-day vulnerabilities are found with over \$2,500 bounty rewards.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 2.2. Running Example
To illustrate the basic idea of how Verite finds profitable vulnerabilities and exploits them, we provide an adapted example from a real-world project VULN1  [^1] Due to double-anonymized review, we can’t reveal the real project names here. where Verite can maximize profits to highlight its key designs in Figure 1. The target contract on the right is vulnerable because the burn function is wrongly declared as public, allowing anyone to burn other users’ tokens, which typically can lead to a price manipulation attack. In the following, we will explain how the attack can be generated by Verite and why existing solutions are limited.
@@ 2.3. Our Approach
Figure 2 illustrates the overall architecture of Verite, an on-chain fuzzing platform with a profit-centric fuzzing loop. Initially, users provide the target contract address, and Verite fetches the code and interfaces of the target contract accordingly and starts fuzzing. During each iteration of fuzzing, Verite generates an attack contract by combining generated actions and accounting oracles, then deploys this attack contract to the EVM. The lower left part of the figure depicts the general fuzzing loop of Verite, which employs an action-based mutator (Section 3.2) to generate transactions with diverse fund flows. Throughout the fuzzing process, if any input seed is identified as potentially profitable (Section 3.3), it is elevated to the SGD-based profit maximizer (Section 3.4) for parameter optimization. The input that ultimately yields the highest positive profit (Section 3.5) is returned as the output of the fuzzer.
@@ 18             $\mathcal{I}\leftarrow I^{\prime}$
Smart contract fuzzing presents a challenging search space where both API sequences and their parameters must be semantically correct to execute transactions effectively. Existing fuzzers (Choi et al., 2021a; Shou et al., 2023) typically generate transactions at the API level by randomly selecting API sequences and parameters, which often proves inefficient and ineffective for producing valid transactions. To improve upon this, Verite employs action-based mutation. In Verite, a transaction $T_{i}$ within input $I$ can either be a call to external contract functions or a high-level DeFi action   [^2] Internally represented as a transaction to the attacker contract itself. composed of a series of predefined sub-transactions. Although new and customized actions can theoretically be introduced continuously, Verite summarizes eight actions that are essential for detecting profitable vulnerabilities in the current smart contract ecosystem. Table 1 provides details on these actions. One of the actions is the transfer functionality of the ERC20 (erc, 2015) token, while the others are common operations supported by leading decentralized exchanges (DEXs) such as Uniswap (Adams et al., 2020, 2021) and DODO (dod, 2024). These actions were chosen due to their significant role in the ecosystem: Uniswap and DODO alone account for over 75% and 90% of the weekly trading volumes of ERC20 tokens on Ethereum mainnet (ETH) and Binance Smart Chain (BSC), respectively (def, 2024b). The actions are implemented by rewriting code snippets from documentation samples or unit tests (uni, 2024). By selecting these prevalent on-chain projects, Verite ensures it targets the most impactful areas for vulnerability detection. Furthermore, to enhance flexibility and support a broader range of actions, Verite automatically detects actions defined in the attacking contract. This feature allows for the extension of actions by writing new ones in Solidity without needing modifications to Verite’s internal fuzzing implementation. More guidance on semi-automated extending supported actions with LLM-based methods can be found on our website (our, 2024).
@@ 3.4. Stochastic Gradient Descent Based Profit Maximizer
Regarding the parameter $p,q,n$, the basic intuition of the Equation 5 is that if the gradient direction in the $j+1$ loop is the same as the previous loop, we will take a bigger step, expecting a faster convergence. With the opposite gradient direction, we will take a reduced step with an opposite direction to avoid diverging from the maximum point. The step of $x_{i}$ from repetitions are separately calculated with smaller steps to avoid gradient vanishing because of revert. The three parameters strongly correlate with properties of the profit function $P$, which doesn’t have an analytic form. Verite empirically uses $p=\frac{3}{2},q=-\frac{1}{3},n=5$ for all experiments of this paper.
@@ 4. Evaluation
Implementation    Verite is written from scratch with 10,106 lines of rust code and 822 lines of solidity code. We use the revm (rev, 2024) as our EVM emulator and libafl (Fioraldi et al., [n. d.]) to build our fuzzer. There are roughly 300 lines of solidity code to define actions and 2,000 lines of Rust code to identify and do action-based mutation automatically. Verite is highly flexible and modularized and easy to migrate to other networks since we abstract over the execution engines and assets accounting. In addition, we also implemented four variants of Verite for the ablation study to answer RQ2:
@@
Baseline    To answer RQ1, we evaluated Verite against two baselines: ItyFuzz (Shou et al., 2023) [^4] Commit hash: 8fdcdf7851c7e66ec8e435d555ef4a4fd217a3b6 and on-chain real-world attacking transactions of the incidents collected in our datasets. ItyFuzz is picked because it is the well-maintained, state-of-art on-chain fuzzer. The real-world attacking transactions are usually highly optimized because hungry on-chain searchers tend to instantly fill any residual profits due to the "Dark Forest" nature of the blockchain(Qin et al., 2022). We further excluded fuzzers that only support finding profitable vulnerabilities in native currency because of the popularity of ERC20 tokens. It will take non-trivial extra development efforts to modify them to support detecting ERC20-related profitable vulnerabilities.
@@
Sum 106,432 51,927,399.2 18,338,315.7 92.2% 94.7% 4.1% 6.7%
FP 17555The 11 targets: valuedefi, pancakehunny, roi, tinu, lw, selltoken, sellc03, cfc, shido, lusd, newfi, are omitted in the table because only ItyFuzz finds False Positives on them and Verite reports nothing. - 0 0 0 0 0
@@
Dataset Construction    To evaluate Verite, we extended the dataset ItyFuzz used (Shou et al., 2023) by another half year following the same criteria, from the same data source, DefiHackLabs (def, 2024a). Our new dataset consists of 61 open-source Web3 projects relating to Uniswap and ERC20 tokens exploited on the Ethereum mainnet (ETH) or Binance Smart Chain (BSC) in the past two years, including a variety of bug types. The loss amount ranges from hundreds to millions of dollars, with an average of \$1.1 million per project. We extracted all relevant participant addresses (except the attacker and attacker-controlled contracts) from the attack and used the block number just before the attack to fork the specific chain. The full dataset can be found on our website (our, 2024).
@@ 4.1. RQ1 State-of-the-Art Comparison
② Analysis on Targets Verite Failed to Exploit    Although Verite proves its value by yielding a large amount of profits with concrete proof-of-profits, there are still a few targets where Verite fails to find any positive profits. We also manually analyze these cases and the attacking transactions. The first reason for the false negatives is the magic numbers or addresses required for attacks to succeed, which are almost impossible to brute-force without further approaches like symbolic execution. Besides, other attacks have rather strong control or data flow dependency, which Verite fails to satisfy during fuzzing. Employing symbolic execution or dynamic dataflow analysis or extracting more actions automatically shall mitigate the limitations.
@@
Compared with ItyFuzz    ItyFuzz can only generate profitable attacks on nine targets, at a sum of 100k profits while Verite outperforms ItyFuzz by finding more profits on all the nine targets and yielding 172 times more total profits.
@@
Imbalanced Uniswap Pairs 37 9
@@
Total 45 9
Table 3. Breakdown of proof-of-profit candidates. ItyFuzz only implemented "Imbalanced Uniswap Pairs". Some targets may satisfy more than 1 rule.
② False Positive Analysis    On-chain arm-race is highly automated, and thus, one of the goals of Verite is to reduce false positives as much as possible because false positives will take extra manual efforts for further validation. Of all the targets we evaluated, Verite has no false positives while ItyFuzz has 17 false positives. The reasons for the false positives of ItyFuzz are illegal or unreachable EVM states because of hijacking EVM execution or error during off-chain accounting emulation [^8] Detailed false positive analysis and proof of ItyFuzz will be available at our website (our, 2024).. For instance, transactions that shall revert could still be committed during ItyFuzz ’s emulation, leading to false positives. In contrast, Verite is equipped with on-chain accounting and ensures the profits gained are concrete and reproducible by leaving EVM execution intact.
③ Throughput and Coverage    We collected the instruction coverage and transaction throughput of Verite and ItyFuzz following previous work (Wu et al., 2024; Shou et al., 2023). Verite averagely achieves 3,178 transactions per second and 66.2% instruction coverage while ItyFuzz only achieves 1,729 transactions per second and 54.0% instruction coverage. ItyFuzz has a lower transaction throughput mainly because of the overhead incurred by taking huge amounts of snapshots due to the complex state space of on-chain fuzzing. We attributed the coverage improvements to our faster throughput and profits-targeted fuzzing strategies. The high throughput and coverage ensure Verite is competitive for on-chain auditing arm-race.
@@ 4.2. RQ2 Ablation Study
On-Chain Accounting    Verite-NOACC behaves much worse than Verite by exploiting 13 less targets. The first reason is that some inputs can have already exploited vulnerabilities, but without accounting, Verite-NOCDT fails to determine they are profitable attacks. In addition, the lack of withdrawal of all assets also affects the SGD process, which heavily relies on the profit function to report profits, as targets like health and bevo suggest.
Proot-of-Profit Candidate and SGD    Verite-NOCDT and Verite-NOGRD have a similar performance by missing seven targets compared with Verite because both variants do not exploit the profit potential of the inputs with negative inputs. To illustrate this process better, we chose two representative cases gss and health to plot how the maximum profits change with the fuzzing campaign going [^9] Figures of all targets can be found on our website (our, 2024).. The gss case is typical in that SGD starts from a proof-of-profit and maximizes the profits until converging. In contrast, health case illustrates how SGD exploits the profit potential of inputs with negative profits. In both graphs, the action-based mutation and SGD contribute to the profit increase alternatively.

### v2

@@ Abstract.
To address these gaps, we have developed Verite, a profit-centric smart contract fuzzing framework that not only effectively detects those profitable vulnerabilities but also maximizes the exploited profits. Verite has three key features: ① DeFi action-based mutators for boosting the exploration of transactions with different fund flows; ② potentially profitable candidates identification criteria, which checks whether the input has caused abnormal fund flow properties during testing; ③ a gradient descent-based profit maximization strategy for these identified candidates. Verite is fully developed from scratch and evaluated on a dataset consisting of 61 exploited real-world DeFi projects with an average of over 1.1 million dollars loss. The results show that Verite can automatically extract more than 18 million dollars in total and is significantly better than state-of-the-art fuzzer ItyFuzz in both detection (29/10) and exploitation (134 times more profits gained on average). Remarkably, in 12 targets, it gains more profits than real-world attacking exploits (1.01 to 11.45 times more). Verite is also applied by auditors in contract auditing, where 6 (5 high severity) zero-day vulnerabilities are found with over \$2,500 bounty rewards.
Smart Contract, Fuzzing, Vulnerabilities Research
Cen Zhang is the corresponding author.
conference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYcopyright: ccdoi: 10.1145/3715720journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE008publicationmonth: 7ccs: Security and privacy Software security engineeringccs: Software and its engineering Search-based software engineeringccs: Security and privacy Vulnerability scanners
@@ 2.2. Running Example
To illustrate the basic idea of how Verite finds profitable vulnerabilities and exploits them, we provide an adapted example from a real-world project ENDERV1 where Verite can maximize profits to highlight its key designs in Figure 1. The target contract on the right is vulnerable because the burn function is wrongly declared as public, allowing anyone to burn other users’ tokens, which typically can lead to a price manipulation attack. In the following, we will explain how the attack can be generated by Verite and why existing solutions are limited.
@@ 2.3. Our Approach
Figure 2 illustrates the overall architecture of Verite, an on-chain fuzzing platform with a profit-centric fuzzing loop. Initially, users provide the target contract address, and Verite fetches the code and interfaces of the target contract accordingly and starts fuzzing. During each iteration of fuzzing, Verite generates several transactions, each of which encodes generated actions and accounting oracles using the multicall pattern [^1] https://github.com/OpenZeppelin/openzeppelin-contracts/blob/v4.9.5/contracts/utils/Multicall.sol, and then calls the attack contract with the transactions. The lower left part of the figure depicts the general fuzzing loop of Verite, which employs an action-based mutator (Section 3.2) to generate transactions with diverse fund flows. Throughout the fuzzing process, if any input seed is identified as potentially profitable (Section 3.3), it is elevated to the SGD-based profit maximizer (Section 3.4) for parameter optimization. The input that ultimately yields the highest positive profit (Section 3.5) is returned as the output of the fuzzer.
@@ 18             $\mathcal{I}\leftarrow I^{\prime}$
Smart contract fuzzing presents a challenging search space where both API sequences and their parameters must be semantically correct to execute transactions effectively. Existing fuzzers (Choi et al., 2021a; Shou et al., 2023) typically generate transactions at the API level by randomly selecting API sequences and parameters, which often proves inefficient and ineffective for producing valid transactions. To improve upon this, Verite employs action-based mutation. In Verite, a transaction $T_{i}$ within input $I$ can either be a call to external contract functions or a high-level DeFi action   [^2] Internally represented as a transaction to the attacker contract itself. composed of a series of predefined sub-transactions. Although new and customized actions can theoretically be introduced continuously, Verite summarizes eight actions that are essential for detecting profitable vulnerabilities in the current smart contract ecosystem. Table 1 provides details on these actions. One of the actions is the transfer functionality of the ERC20 (erc, 2015) token, while the others are common operations supported by leading decentralized exchanges (DEXs) such as Uniswap (Adams et al., 2020, 2021) and DODO (dod, 2024). These actions were chosen due to their significant role in the ecosystem: Uniswap and DODO alone account for over 75% and 90% of the weekly trading volumes of ERC20 tokens on Ethereum mainnet (ETH) and Binance Smart Chain (BSC), respectively (def, 2024b). The actions are implemented by rewriting code snippets from documentation samples or unit tests (uni, 2024). By selecting these prevalent on-chain projects, Verite ensures it targets the most impactful areas for vulnerability detection. Furthermore, to enhance flexibility and support a broader range of actions, Verite automatically detects actions defined in the attacking contract. This feature allows for the extension of actions by writing new ones in Solidity without needing modifications to Verite’s internal fuzzing implementation. More guidance on semi-automated extending supported actions with LLM-based methods can be found on our website (our, 2025).
@@
For simplicity, Verite does not make any explicit assumptions about the inter-action dependencies, meaning any two actions can be combined freely during mutation.
@@ 3.4. Stochastic Gradient Descent Based Profit Maximizer
Regarding the parameter $p,q,n$, the basic intuition of the Equation 5 is that if the gradient direction in the $j+1$ loop is the same as the previous loop, we will take a bigger step, expecting a faster convergence. With the opposite gradient direction, we will take a reduced step with an opposite direction to avoid diverging from the maximum point. The step of $x_{i}$ from repetitions are separately calculated with smaller steps to avoid gradient vanishing because of revert. The three parameters strongly correlate with properties of the profit function $P$, which doesn’t have an analytic form. Therefore, there is no general optimal step parameter. However, as long as the parameters remain in reasonable values, i.e., $p-1$ is not too close to 0, $p$ is not excessively large, $p>-q$ and $n\leq 9$, the impact of parameters’ selection is quite limitted because we actually adopt a variant of binary search with lookback mechanism. Verite empirically uses $p=\frac{3}{2},q=-\frac{1}{3},n=5$ for all experiments of this paper.
@@ 4. Evaluation
Implementation    Verite is written from scratch with 10,106 lines of Rust code and 822 lines of Solidity code. We use the revm (rev, 2024) as our EVM emulator and libafl (Fioraldi et al., 2022) to build our fuzzer. There are roughly 300 lines of Solidity code to define actions and 2,000 lines of Rust code to identify and do action-based mutation automatically. Verite is highly flexible and modularized and easy to migrate to other networks since we abstract over the execution engines and assets accounting. In addition, we also implemented four variants of Verite for the ablation study to answer RQ2:
@@
Baseline    To answer RQ1, we evaluated Verite against two baselines: ItyFuzz  (Shou et al., 2023) [^4] Commit hash: 8fdcdf7851c7e66ec8e435d555ef4a4fd217a3b6 and on-chain real-world attacking transactions of the incidents collected in our datasets. ItyFuzz is picked because it is the well-maintained, state-of-art on-chain fuzzer. The real-world attacking transactions are usually highly optimized because hungry on-chain searchers tend to instantly fill any residual profits due to the "Dark Forest" nature of the blockchain (Qin et al., 2022). We further excluded fuzzers that only support finding profitable vulnerabilities in native currency because of the popularity of ERC20 tokens. It will take non-trivial extra development efforts to modify them to support detecting ERC20-related profitable vulnerabilities. Specifically, we once evaluated ConFuzzius (Torres et al., 2021) and EF/CF (Rodler et al., 2023) with the same experiment setup on our dataset. However, aligning with recent research work (Zhang et al., 2023), none of them found any proof-of-profit due to lacking concepts modeling like tokens and liquidty, which our on-chain profit accounting oracle does.
@@
Sum 106,474 51,927,399.2 18,338,315.7 92.2% 94.7% 4.1% 6.7%
FP 16555The 11 targets: valuedefi, pancakehunny, roi, tinu, lw, selltoken, sellc03, cfc, shido, lusd, newfi, are omitted in the table because only ItyFuzz finds False Positives on them and Verite reports nothing. - 0 0 0 0 0
@@
Dataset Construction    To evaluate Verite, we extended the dataset ItyFuzz used (Shou et al., 2023) by another half year following the same criteria, from the same data source, DefiHackLabs (def, 2024a). Our new dataset consists of 61 open-source Web3 projects relating to Uniswap and ERC20 tokens exploited on the Ethereum mainnet (ETH) or Binance Smart Chain (BSC) in the past two years, including a variety of bug types. The loss amount ranges from hundreds to millions of dollars, with an average of \$1.1 million per project. We extracted all relevant participant addresses (except the attacker and attacker-controlled contracts) from the attack and used the block number just before the attack to fork the specific chain. The full dataset can be found on our website (our, 2025).
@@ 4.1. RQ1 State-of-the-Art Comparison
② Analysis on Targets Verite Failed to Exploit    Although Verite proves its value by yielding a large amount of profits with concrete proof-of-profits, there are still a few targets where Verite fails to find any positive profits. We also manually analyze these cases and the attacking transactions to understand the failure reasons.
  • Hard Constraints The first reason for the false negatives is the hard constraints, including magic numbers, addresses, ABI decoding for bytes and block or timestamp constraints. 17 cases fall in this category. Employing symbolic execution or dynamic dataflow analysis or extracting more actions automatically shall mitigate the limitations.
  • Inter-Action Dependency For instance, the staking pool related functions expect the previous returned staking ID as its argument. This includes 6 cases.
  • Limited Computation Resources Though theoretically solvable, Verite fails to exploit the left nine targets within the given CPU cores and fuzzing duration.
@@
Compared with ItyFuzz    ItyFuzz can only generate profitable attacks on ten targets, at a sum of 100k profits while Verite outperforms ItyFuzz by finding more profits on all the ten targets and yielding 172 times more total profits.
@@
Imbalanced Uniswap Pairs 37 10
@@
Positive Profits 22 10
Total 45 19
Table 3. Breakdown of proof-of-profit candidates. ItyFuzz only implemented "Imbalanced Uniswap Pairs" and "Positive Profits". Some targets may satisfy more than 1 rule.
② False Positive Analysis    On-chain arm-race is highly automated, and thus, one of the goals of Verite is to reduce false positives as much as possible because false positives will take extra manual efforts for further validation. Of all the targets we evaluated, Verite has no false positives while ItyFuzz has 16 false positives. The reasons for the false positives of ItyFuzz are illegal or unreachable EVM states because of hijacking EVM execution or error during off-chain accounting emulation [^8] Detailed false positive analysis of ItyFuzz will be available at our website (our, 2025).. For instance, transactions that shall revert could still be committed during ItyFuzz ’s emulation, leading to false positives. In contrast, Verite is equipped with on-chain accounting and ensures the profits gained are concrete and reproducible by leaving EVM execution intact.
1function _transfer(address from, address to, uint256 amount ) internal {
2 if (to == address(pair) from == address(pair) ) {
3 if (takeFee && !exclusiveFromFee[from]) {
4 fee = amount.mul(rate).div(1000);
5 _balance[from] = _balance[from].sub(amount).sub(fee);
6 _balance[destroyAddress] = _balance[destroyAddress].add(fee);
7 emit Transfer(from, destroyAddress, fee);
8 }
9 }
10}
Listing 4: The code snippet of critical lines of dfs
③ Throughput and Coverage    We collected the instruction coverage and transaction throughput of Verite and ItyFuzz following previous work (Wu et al., 2024; Shou et al., 2023). Verite averagely achieves 3,178 transactions per second and 66.2% instruction coverage while ItyFuzz only achieves 1,729 transactions per second and 54.0% instruction coverage. 50 out of 61 projects reach the maximum coverage in first 6 hours while the rest in 12 hours. ItyFuzz has a lower transaction throughput mainly because of the overhead incurred by taking huge amounts of snapshots due to the complex state space of on-chain fuzzing. We also fetched the source code mappings from block explorers and analyze the line coverage of the targets that only Verite finds proof-of-profits. It turns out that 8 out of 14 targets [^9] Six targets do not have source mapping. have higher coverage and reach the lines ItyFuzz fails to cover due to our action-based mutation and profits-targeted fuzzing strategies. For instance, We use dfs [^10] Address: 0x2B806e6D78D8111dd09C58943B9855910baDe005 as a case study as shown in Listing 4. dfs takes a fee on selling tokens but asks users to calculate the fee in advance, i.e. users can not sell 100 tokens even if users hold exactly 100 tokens. Our action-based mutation supports mutating the ratio of transferring and selling tokens, as detailed in the Section 3.2. Thus, we can have higher code coverage by finishing the swapping successfully while ItyFuzz always reverts at Line 5. The high throughput and coverage ensure Verite is competitive for on-chain auditing arm-race.
@@ 4.2. RQ2 Ablation Study
On-Chain Accounting    Verite-NOACC behaves much worse than Verite by exploiting 13 less targets. The first reason is that some inputs could have already exploited vulnerabilities, but without accounting, Verite-NOCDT fails to determine they are profitable attacks. In addition, the lack of withdrawal of all assets also affects the SGD process, which heavily relies on the profit function to report profits, as targets like health and bevo suggest.
Proot-of-Profit Candidate and SGD    Verite-NOCDT and Verite-NOGRD have a similar performance by missing seven targets compared with Verite because both variants do not exploit the profit potential of the inputs with negative inputs. To illustrate this process better, we chose two representative cases gss and health from RQ1 to plot how the maximum profits change with the fuzzing campaign going in Figure 3. The gss case is typical in that SGD starts from a proof-of-profit and maximizes the profits until converging. In contrast, health case illustrates how SGD exploits the profit potential of inputs with negative profits. In both graphs, the action-based mutation and SGD contribute to the profit increase alternatively. On average, the time distribution of the action-based mutation exploration stage and SGD stage is 52:48 with the specific ratio depending on the number of proof-of-profit candidates. For projects like gss and health where proof-of-profits candidates are continously found, Verite could spend siginificant amount of time on the SGD stage while for projects with few or no proof-of-profit canditates, SGD stage takes much less time. For both stages, EVM emulation accounts for almost all ($\geq 99\%$) computation resources.  [^11] Figures of the time distributions, profiles and profit increasing contributions of all targets can be found on our website (our, 2025).